OpenStudy (anonymous):

The area of a rectangular pool is given by the trinomial 8y2 + 4y – 12. What are the possible dimensions of the pool? Use factoring.

OpenStudy (arnavguddu):

8y2+4y-12 = 4(2y2+y-3) = 4(2y2+3y-2y-3)=4(y(2y+3)-(2y+3))=4(y-1)(2y+3)=0 say so either y-1=0 or 2y+3=0 or y=1 or y=-3/2 as dimension cannot be negative.so valid y for rectangle is 1

OpenStudy (hoblos):

8y2 + 4y – 12 =2(4y2 + 2y -6) = 2(2y-2)(2y+3) L=4y+6 W=2y-2 OR L= 2y+3 W=4y-4

OpenStudy (anonymous):

(–2y – 2) and (–4y – 6) (2y – 2) and (4y + 6) (2y – 2) and (4y + 6) (2y + 2) and (4y – 6)

OpenStudy (hoblos):

from the options (2y – 2) and (4y + 6) fits